OpenStudy (anonymous):

If f(x)=3x+5 and g(x)=x-1, then f[g(0)]=?

OpenStudy (anonymous):

\[f(g(x))=f(x-1)=3(x-1)+5~~\Rightarrow~~f(g(0))=?\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

thx